ST. Calculate the variance of a random variable ξ, such that P(ξ=0)=0.2,P(ξ=1)=0.3,P(ξ=2)=0.5.


1
Expert's answer
2021-12-27T04:20:15-0500
"E(\xi)=0.2(0)+0.3(1)+0.5(2)=1.3"

"E(\xi^2)=0.2(0)^2+0.3(1)^2+0.5(2)^2=2.3"

"Var(\xi)=E(\xi^2)-(E(\xi))^2"

"=2.3-(1.3)^2=0.61"

"Var(\xi)=0.61"